OpenStudy (anonymous):

You and two friends visit a pizza shop. There are 13 toppings. 8 toppings are meat and the rest are vegtables. Each of you orders a different topping(no repeats) at random. What is the probability that your group orders only meat toppi

OpenStudy (alekos):

P(meat) = 8/13 P(group orders meat) = P(1st orders meat) x P(2nd orders meet) x p(3rd orders meat)